Consider the following  information:  

Marginal propensity to consume = 0,9

Investment = R180 million

Autonomous consumption = R105 million. The equilibrium level of income is:-

Question 2 options:



1
Expert's answer
2021-12-12T16:31:56-0500

At Equillibrium, AD=AS

Y=C+I

C=a+bY

C=R105+0.9Y

Y=R105+0.9Y+R108

Y-0.9Y=R213

0.1Y=R213

Y=R213/0.1

=R2130